OpenStudy (anonymous):

f(x)=x^2 - x- 6 determine the x- intercepts of the quadratic function i think it (2,0)(-3,0)

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

To get x - intercepts you let y = 0 (in this case f(x) = 0).

OpenStudy (anonymous):

x^2 +x - 6 = 0 x^2 +3x -2x -6 = 0 x(x+3) -2(x+3) = 0 (X-2)(x+3)= 0 answer = (2,0)(-3,0)

OpenStudy (anonymous):

that how i got the answer did i do it right?

OpenStudy (anonymous):

(x-3)(x+2)

OpenStudy (anonymous):

ok i see what did wrong thank you for all your help

OpenStudy (anonymous):

np